Circular No. 6143

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

JUPITER I (IO) J. R. Spencer, J. E. B. Spencer, and D. Griep, NASA Infrared Tele-

scope Facility (IRTF), report their observation at 3.5–4.8 μm of a major volcanic outburst on Io, observed during Mar. 2.66 to 2.75 UT: “The outburst was seen in disk-resolved images at the IRTF and is also apparent in preliminary 5-μm photometry. The location of the spot in the images is very roughly near longitude 130^{o}, latitude −20^{o} (visual estimate only), so it is still rotating onto the disk at this time. Such outbursts tend to last only a few hr, so further imaging, photometry, or spectroscopy is requested.”

POSSIBLE NOVA IN THE LARGE MAGELLANIC CLOUD W. Liller, Instituto Isaac Newton, Viña del Mar, reports his discovery

of a possible nova on a pair of Technical Pan films taken of the LMC on Mar. 2.11 UT. The object is located at α = 5^{h}26^{m}.8, δ = −70^{o}01′ (equinox 2000.0); this is ∼ 3′ northeast of a star of mag 7.9. Approximate magnitude estimates for the possible nova: Feb. 25.146, 12.6; Mar. 2.11, 10.7.

X-RAY NOVA 1994 IN SCORPIUS N. Alexandrovich, K. Borozdin, and R. Sunyaev, on behalf of the MIR-

Kvant team, report: “The superluminal radio source GRO J1655–40 (cf. IAUC 6046, 6051, 6056, 6078, 6087, etc.) was found to be very bright during a TTM observation on Feb. 22. The flux in the energy band 3–5 keV was close to 1.7 Crab. The source spectrum was soft and could be approximated by a blackbody fit, with temperature kT = 1.04 ± 0.02 keV in the energy band 2–12 keV.”

1995 DA_{2}, 1995 DB_{2}, 1995 DC_{2}

MPEC 1995–E05, 1995–E06, and 1995–E07 contain reports of the discovery by J. Luu (Center for Astrophysics) and D. Jewitt (University of Hawaii) of three new transneptunian objects. Circular orbits computed by B. G. Marsden have radii of 34.283, 40.953, and 45.406 AU, respectively. 1995 DA_{2} and DB_{2} are almost opposite Neptune and within ∼ 1^{o} of each other. Marsden notes that the stability of 1995 DA_{2} is best assured by supposing this object to be near perihelion in a 1:2 Neptune-librating orbit, and he suggests that 1995 DB_{2} may represent a similar—but less extreme—case.

1995 March 2 Daniel W. E. Green